OpenStudy (karamarie02):

The answer is B I took the test Which answer best describes the solution to the equation 3m – 9 = 9 + 3m? A. one solution B. no solution C. an infinite number of solutions D. m = 27

OpenStudy (texaschic101):

u tried solving equation yet ?

OpenStudy (cirenawilson):

C, can i get a medal?

OpenStudy (texaschic101):

3m - 9 = 9 + 3m....subtract 3m from both sides 3m -3m - 9 = 9 + 3m - 3m -9 = 9....incorrect...no solution

OpenStudy (karamarie02):

The answer was B

OpenStudy (shansha1):

3m - 9 = 9 + 3m....subtract 3m from both sides 3m -3m - 9 = 9 + 3m - 3m -9 = 9....incorrect...no solution